A paper airplane follows a parabolic path with h=-1/4t^2+t+3, where h is the height in meters, and t is the time in seconds. Algebraically determine how long it takes for the paper airplane to hit the ground.
Answer by ankor@dixie-net.com(22740)

A paper airplane follows a parabolic path with h=-1/4t^2+t+3, where h is the height in meters, and t is the time in seconds. Algebraically determine how long it takes for the paper airplane to hit the ground.
:
When the plane hits the ground, h = 0. so we have:
-1/4t^2 + t + 3 = 0
multiply equation by -4 to get rid of the denominator & changes the sign.
t^2 - 4t - 12 = 0
factor
(t - 6)(t + 2) = 0
Positive solution
t = 6 seconds
:
:
Check solution in original equation: t=6
h = -1/4(6^2) + 6 + 3
h = -1/4(36) + 6 + 3
h = -9 + 9
h = 0